The area of guardriateral formed by the lines y = 2x + 3, y = 0, x = 4, x = 6 is equal to which of the following ?(A) 26 square unit (B) 24 square unit (C) 20 square unit (D) None of these

Grade:12

2 Answers

venkateshwarrao
7 years ago
Area of quadrilateral=area of rectangle ABCD+ area of triangle CDE=AB*AD+1/2*CD*CE=2*11+1/2*2*4=22+4=26THE. Line y=2x+3 meets x,y axis at(-3/2,0)&(0,3)and extended it , this line will meet the line sx=4&x=6at D(4,11)&E(6,15) where A(4,0)&B(6,0) on x axis,draw perpendicular from D to BE we get C(6,11), therefore ABCD is rectangle&CDE IS a triangle so area is 26 sq units "OK": all the best
venkateshwarrao
7 years ago
Sorry,the answer,is 26-2=24.because the side CD is common to rectangle and triangle,therefour area of quadrilateral is 24 sq units